OpenStudy (reneeshumpert):

7x - 3y = 4 2x - 4y = 1 Which of the following systems is not the same as solving the system shown? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14 .. am I supposed to be multiplying them?

OpenStudy (reneeshumpert):

the equations

OpenStudy (texaschic101):

look at your first equation...7x - 3y = 4 now what would you have to multiply 7x by to get 28x.....you would have to multiply by 4. so 4 * (7x - 3y = 4) = 28x - 12y = 16 now check out your second equation..2x - 4y = 1 now what would you have to multiply 2x by to get -6x....that would be -3 so -3 (2x - 4y = 1) = -6x + 12y = -3 so I would say the first one works.....do you see what I did ? Do you understand ? If they didn't match, I would have gone to the other answer choices

OpenStudy (reneeshumpert):

No I'm not really understand , explain a little more? And future reference I like to walk through the problem lol not just get the answer.

OpenStudy (reneeshumpert):

like where did the 4 come in at ? or that was an educated guess?

OpenStudy (texaschic101):

now if I wanted to try the second answer choice....7x - 3y = 4 what would you have to multiply 7x by to get 14x....you would have to multiply by 2 2(7x - 3y = 4) = 14x - 6y = 8.....that does not match...not the answer

OpenStudy (reneeshumpert):

Oh so basically you're using the answer choices to fill in for the equation ?

OpenStudy (texaschic101):

I am basically multiplying the original equations by a certain number to see if it matches the answer choices....does that make sense ??

OpenStudy (reneeshumpert):

Oooo yeah that makes sense, I get it now

OpenStudy (reneeshumpert):

but how did you come up with those numbers?

OpenStudy (texaschic101):

ok...your original equation....7x - 3y = 4 I was just trying the first answer choice...28x + 12y = 16... the number I want to multiply by was found by 7x * what = 28x.....so the number I want to multiply the entire equation by is 4 4(7x - 3y = 4) 28x - 12y = 16 Is this making any sense ?

OpenStudy (reneeshumpert):

your answer was actually incorrect. thank you tho

OpenStudy (texaschic101):

oh..I am so sorry....your right... 7x - 3y = 4...multiply by -4 -4(7x - 3y = 4) = -28x + 12y = -16 so sorry...answer is the third one